### v1

@@ Abstract
In this work-in-progress, we investigate the certification of artificial intelligence (AI) systems, focusing on the practical application and limitations of existing certification catalogues by attempting to certify a publicly available AI system. We aim to evaluate how well current approaches work to effectively certify an AI system, and how publicly accessible AI systems, that might not be actively maintained or initially intended for certification, can be selected and used for a sample certification process. Our methodology involves leveraging the Fraunhofer AI Assessment Catalogue as a comprehensive tool to systematically assess an AI model’s compliance with certification standards. We find that while the catalogue effectively structures the evaluation process, it can also be cumbersome and time-consuming to use. We observe the limitations of an AI system that has no active development team anymore and highlighted the importance of complete system documentation. Finally, we identify some limitations of the certification catalogues used and proposed ideas on how to streamline the certification process.
@@ 1 Introduction
Artificial intelligence (AI) has evolved, over several decades, to complex machine learning (ML) algorithms and neural networks that are common-place today. However, in recent years, AI systems were increasingly integrated into our daily lives, moving from specialized research labs to mainstream applications [1]. Today AI is utilized in critical fields such as healthcare [2], human resources [3], social networks and recommender systems [4, 5], as well as finance [6]. AI now plays a significant role in shaping our interactions with technology and informing decision-making processes across various sectors [7]. The rapid proliferation of AI applications has raised concerns about safety, privacy, fairness, and further ethical implications [8, 9]. In response to these challenges, AI governance has become an increasingly prominent focus for legislators and policymakers worldwide [10]. The European Union’s AI Act, for instance, represents a landmark piece of legislation that aims to establish a comprehensive regulatory framework for AI systems [11]. Similar initiatives are underway in other countries and regions. This reflects a growing global consensus on the need for AI governance [12]. Organizations and policymakers are increasingly emphasizing the need for AI system certification to achieve regulatory compliance and foster confidence in AI systems. Since the EU AI Act entered into force on August 1, 2024, this is now more important than ever. The certification of AI systems is a complex challenge, distinctly different from traditional software certification. While conventional software certification primarily focuses on functionality and security, AI certification must address a broader spectrum of concerns, including prediction accuracy, fairness, transparency, and other ethical considerations [13].
@@ 2 Current State of AI Regulation
Rapid advancement and widespread adoption of AI in various industries require the development of comprehensive regulatory frameworks, and increasing AI deployment in more critical areas appears to require stricter regulation [15]. Some bodies like the US Food and Drug Administration have already approved certain AI applications, particularly in medicine [16]. In the European Union, some medical applications also got approval, one example is the ChestLink software, that automatically reports chest x-rays that are classified as normal. Systems such as this set an important precedent for other medical and potentially nonmedical systems in the future [17]. Recently, lawmakers have recognized the urgent need for comprehensive regulation. New regulatory frameworks intend to encompass not only individual industries, but also to create universal rules that ensure consistency and fairness. Worldwide, several organizations are advancing policy initiatives [12]. In the following, we describe some key efforts in this area.
@@ 2.1 EU Artificial Intelligence Act
The European Union’s AI Act is currently the most significant and far-reaching regulatory initiative in the field of AI, and it took effect on 1st of August 2024. The impact of the AI Act will extend far beyond the EU’s boarders, potentially setting global standards for AI management. Key Objectives of the AI Act are: [18]:
@@
The AI Act uses a risk-based approach to achieve these goals.
@@ 3 Certifying AI
Certifying AI systems also presents unique challenges due to their complex and often opaque nature. Often, humans do not directly program logical rules to model decision-making processes. Instead, AI systems use various methods to analyse and interpret data, learning their own rules for decision-making processes. Humans do have multiple ways to influence and understand how the system makes decisions and how good the outcome is. However, the process lacks complete transparency. This complexity necessitates a different approach to certification, emphasizing the need to establish a comprehensive framework that accounts for the inherent opacity and adaptability of AI technologies [13]. Moreover, AI systems can evolve and change their behaviour over time by retraining with new data. This continuous learning process adds another layer of complexity to certification, as regulators and developers must continuously monitor and re-evaluate systems to ensure they comply with ethical standards and avoid biases. Addressing these dynamic aspects is essential to develop robust certification practices that can keep pace with the rapidly evolving nature of AI [26].
@@ 4.1 Our Decision to Use the EmoPy Framework and RIOT Project for AI Certification
After careful consideration, we decided to use the EmoPy Framework [29] and its implementation within the RIOT Project [30] for this sample certification. Several key factors influenced this choice, to ensure that both the framework and the project are suitable for the certification process.
@@
Firstly, the EU AI Act’s relevance significantly influenced our decision. Emotion recognition, the primary focus of the EmoPy Framework, aligns with the potential coverage of the EU AI Act. While open-source models like EmoPy are generally exempt, this system may fall under the AI Act’s scope if it was brought to market as a high-risk AI system, a prohibited AI system or an AI system with special transparency obligations.
@@
Another critical factor was the open-source nature and transparency of EmoPy. EmoPy is fully open-source, enabling an in-depth look into the technical details and making the system fully transparent, which should make the certification process possible. The codebase of EmoPy is relatively small and manageable, making it easier to understand and verify. Yet, it is sufficiently large to make a certification worthwhile. Additionally, the framework includes thorough documentation, with multiple articles and resources that describe the model selection process. This level of documentation is critical for reproducibility attempts of the AI models [31], and with this also for the certification process, as it provides clear insights into the design and functionality of the model, facilitating a thorough evaluation. While solid documentation is essential for any certification, it is especially critical for our sample certification, since no company or active development team is managing the project anymore. Although no active development team provides ongoing support, contacting authors and lead developers of the EmoPy framework articles and code, proved beneficial. They kindly addressed questions about the framework and the RIOT setup. Beyond this input, we needed to extract all necessary information from the provided documentation. This reliance on static resources poses a limitation compared to standard certification processes, where ongoing interactions with active developers should be possible.
From a technical perspective, we considered the suitability of the EmoPy framework for sample certification a key factor. Its technical characteristics and comprehensive documentation make it well-suited for the certification process. Additionally, the integration of EmoPy within the RIOT project provides a complete system context, which is essential for certification. Traditional certification catalogues often struggle to validate standalone ML models. However, the RIOT project offers a comprehensive framework where the AI component is embedded within a broader system. This integration is critical, as certification typically applies to an entire system rather than individual components.
@@
All these factors influenced our choice to use the EmoPy framework and the RIOT project for our sample certification. Key considerations included its potential relevance to the EU AI Act, its open-source nature ensuring transparency, and its suitability for a sample certification. Additionally, the comprehensive system context that the RIOT project provides reinforced this decision, enabling an effective and thorough certification process. It is also important to note that both EmoPy and RIOT were not originally intended for certification by their creators. However, they serve well for this academic exercise, as the focus is on the certification procedure rather than the system’s quality or real-world applicability.
@@ 4.2 Brief Overview of the EmoPy Framework
The EmoPy framework provides multiple neural network architectures for facial expression recognition, including ConvolutionalNN, TransferLearningNN, and ConvolutionalLstmNN. These architectures vary in complexity, with the ConvolutionalNN being the most simple, and TransferLearningNN (using Google’s Inception-v3) being the most complex. The authors experimented with different architectures and found that the ConvolutionalNN provides the best overall performance [32]. The EmoPy documentation suggests using two publicly available datasets for training and evaluation: the Microsoft FER2013 dataset and the Extended Cohn-Kanade dataset. The FER2013 dataset contains over 35,000 facial expression images across 7 emotion classes, while the Cohn-Kanade dataset includes 327 facial expression sequences. To increase the size and suitability of the training and validation datasets, the authors applied data augmentation techniques [29].
In EmoPy, developers train the neural networks using a training set and evaluate them on a separate validation set. The process begins by splitting the dataset into training and validation subsets. During training, the network weights are iteratively adjusted to minimize the loss between predicted and labeled emotions. To mitigate overfitting, the authors monitored the gap between training and validation accuracy. This approach prevents overfitting and ensures the model generalizes well by using unseen validation data during training. They measured performance using training and validation accuracy, and analyzed confusion matrices, which visually represent misclassification rates, to help refine the models. Additionally, they performed cross-validation with multiple datasets to confirm the model’s generalizability [29].
The authors tested the neural network architectures on various emotion classification tasks and reported their performance. They identified the ConvolutionalNN model as the best performer, achieving over 90 % accuracy on some emotion subsets. To refine the models further, they analysed confusion matrices to gain insights [32]. The key goals of the EmoPy project are to provide free, open-source, and easy-to-use facial expression recognition capabilities, and to advance research in this field by making the models and datasets publicly available.
@@ 4.3 Brief Overview of the RIOT Project
Figure 1: RIOT Installation in New York 2018. This image shows the RIOT Art installation in New York City in 2018. One participant is standing in front of the screen taking part in the experience [32].
4.4 Summary of the AI Application to be Certified
For this sample certification, the RIOT installation provides the complete system context for the certification process. The EmoPy framework, which uses ML to detect emotions, serves as the core system within this installation. Our certification process focuses on validating the AI system within the RIOT context. The surrounding code, setup, and information that comprise the entire art installation are relevant, as we cannot certify the AI system independently of these components. However, this approach has shortcomings, as not all the required information is available. Our approach allows for an exploration of the certification process while acknowledging its limitations and academic nature. To facilitate the certification process, we prepared a complete overview of all available information about the AI application, which is summarized below. In the following sections, we will delve into the specific certification procedures applied to this facial emotion recognition system.
{mdframed}
[linewidth=1pt]
@@ 4.4 Summary of the AI Application to be Certified
To make the AI certification process feasible, we adhered to the following assumptions:
@@ 5.4 Protection Requirement Analysis
The protection requirement analysis serves as an important first step in the certification process, identifying the risk dimensions that require more in-depth analysis. This analysis involves evaluating the potential impact of the AI system across various dimensions such as fairness, reliability, and data protection. We examined all dimensions and identified several with medium or high risk. For the purposes of this work, we selected two of the required risk dimensions, namely reliability and fairness, for detailed analysis. This selection helps us to focus the sample certification and manage the scope of the certification process. Exploring these two dimensions is also sufficient to understand the certification procedure and draw the appropriate conclusions.
@@ 5.5 Risk Analysis
The risk analysis forms the core of the certification process. We carried out this step by working through a questionnaire from the Fraunhofer Catalogue. The questions addressed different aspects of the selected risk dimensions. For each dimension, we covered topics such as data quality, model design, testing procedures, and operational considerations. Following the individual dimension analyses, we conducted a cross-dimensional assessment to identify potential trade-offs or interactions between the examined dimensions. This step is crucial to ensure a complete understanding of the AI system’s performance and risks.
In the following section, we descuss in detail the challenges we encountered during this certification process. Based on these experiences, we draw conclusions. Additionally, we discuss comparisons with two other certification catalogues to provide a broader perspective on AI certification methodologies.
@@ 7 Conclusion and Future Research Directions
This work-in-progress presents our ongoing research into the practical application of AI certification catalogues, providing insights into both the capabilities and limitations of current certification approaches, as well as the challenges we encountered in certifying an open-source AI system. The implementation of the Fraunhofer AI Assessment Catalogue demonstrated its effectiveness as a comprehensive certification tool, particularly in its systematic approach to evaluate AI systems. We also found that the approach is, at times, bulky and time-consuming in some areas. In future works, other approaches and certification catalogues should be considered, to potentially streamline the process. The certification process highlights the critical importance of complete system documentation and active engagement from the development team. The absence of these elements can negatively impact the certification process and even make certification infeasible. During the certification attempt, we identified the key characteristics an AI system must possess to be adequately certified. Before selecting a public AI system for a sample certification, auditors should consider factors like thorough documentation and accessibility to the development team. These findings directly address the initial research objectives by identifying both useful aspects of this certification approach and areas for improvement. They also provide practical insights into the certification process and its limitations. Future work in this field could focus on developing more flexible certification methodologies that accommodate various system states and development scenarios.
Acknowledgements: This work was supported by the FFG COMET program.

### v2

@@ Abstract
In this work-in-progress, we investigate the certification of AI systems, focusing on the practical application and limitations of existing certification catalogues in the light of the AI Act by attempting to certify a publicly available AI system. We aim to evaluate how well current approaches work to effectively certify an AI system, and how publicly accessible AI systems, that might not be actively maintained or initially intended for certification, can be selected and used for a sample certification process. Our methodology involves leveraging the Fraunhofer AI Assessment Catalogue as a comprehensive tool to systematically assess an AI model’s compliance with certification standards. We find that while the catalogue effectively structures the evaluation process, it can also be cumbersome and time-consuming to use. We observe the limitations of an AI system that has no active development team anymore and highlighted the importance of complete system documentation. Finally, we identify some limitations of the certification catalogues used and proposed ideas on how to streamline the certification process.
@@ 1 Introduction
Artificial intelligence (AI) has evolved, over several decades, to complex machine learning (ML) algorithms and neural networks that are common-place today. However, in recent years, AI systems were increasingly integrated into our daily lives, moving from specialized research labs to mainstream applications [1]. Today AI is utilized in many domains such as healthcare [2], human resources [3], finance [4] as well as in prominent technologies such as recommender systems for e.g., social networks [5, 6] or online marketplaces [7, 8]. AI now plays a significant role in shaping our interactions with technology and informing decision-making processes across various sectors [9]. The rapid proliferation of AI applications has raised concerns about safety, privacy, fairness, and further ethical implications [10, 11]. In response to these challenges, AI governance has become an increasingly prominent focus for legislators and policymakers worldwide [12]. The European Union’s AI Act, for instance, represents a landmark piece of legislation that aims to establish a comprehensive regulatory framework for AI systems [13]. Similar initiatives are underway in other countries and regions. This reflects a growing global consensus on the need for AI governance [14]. Organizations and policymakers are increasingly emphasizing the need for AI system certification to achieve regulatory compliance and foster confidence in AI systems. Since the EU AI Act entered into force on August 1, 2024, this is now more important than ever. The certification of AI systems is a complex challenge, distinctly different from traditional software certification. While conventional software certification primarily focuses on functionality and security, AI certification must address a broader spectrum of concerns, including prediction accuracy, fairness, transparency, and other ethical considerations [15].
@@ 2 Current State of AI Regulation
Rapid advancement and widespread adoption of AI in various industries require the development of comprehensive regulatory frameworks, and increasing AI deployment in more critical areas appears to require stricter regulation [17]. Some bodies like the US Food and Drug Administration have already approved certain AI applications, particularly in medicine [18]. In the European Union, some medical applications also got approval, one example is the ChestLink software, that automatically reports chest x-rays that are classified as normal. Systems such as this set an important precedent for other medical and potentially nonmedical systems in the future [19]. Recently, lawmakers have recognized the urgent need for comprehensive regulation. New regulatory frameworks intend to encompass not only individual industries, but also to create universal rules that ensure consistency and fairness. Worldwide, several organizations are advancing policy initiatives [14]. Now, we describe some key efforts in this area.
@@ 2.1 EU Artificial Intelligence Act
The European Union’s AI Act is currently the most significant and far-reaching regulatory initiative in the field of AI, and it took effect on 1st of August 2024. The impact of the AI Act will extend far beyond the EU’s boarders, potentially setting global standards for AI management. The AI Act uses a risk-based approach to achieve the following goals [20]:
@@ 3 Certifying AI
Certifying AI systems also presents unique challenges due to their complex and often opaque nature. Often, humans do not directly program logical rules to model decision-making processes. Instead, AI systems use various methods to analyse and interpret data, learning their own rules for decision-making processes. Humans do have multiple ways to influence and understand how the system makes decisions and how good the outcome is. However, the process lacks complete transparency. This complexity necessitates a different approach to certification, emphasizing the need to establish a comprehensive framework that accounts for the inherent opacity and adaptability of AI technologies [15]. Moreover, AI systems can evolve and change their behaviour over time by retraining with new data. This continuous learning process adds another layer of complexity to certification, as regulators and developers must continuously monitor and re-evaluate systems to ensure they comply with ethical standards and avoid biases. Addressing these dynamic aspects is essential to develop robust certification practices that can keep pace with the rapid evoluation of AI [28].
@@ 4.1 Our Decision to Use the EmoPy Framework and RIOT Project for AI Certification
After careful consideration, we decided to use the EmoPy Framework [31] and its implementation within the RIOT Project [32] for this sample certification. Several key factors influenced this choice, to ensure that both the framework and the project are suitable for the certification process. Firstly, the EU AI Act’s relevance significantly influenced our decision. Emotion recognition, the primary focus of the EmoPy Framework, aligns with the potential coverage of the EU AI Act. While open-source models like EmoPy are generally exempt, this system may fall under the AI Act’s scope if it was brought to market as a high-risk AI system, a prohibited AI system or an AI system with special transparency obligations.
@@
Another critical factor was the open-source nature and transparency of EmoPy. EmoPy is fully open-source, enabling an in-depth look into the technical details and making the system fully transparent, which should make the certification process possible. The codebase of EmoPy is relatively small and manageable, making it easier to understand and verify. Yet, it is sufficiently large to make a certification worthwhile. Additionally, the framework includes thorough documentation, with multiple articles and resources that describe the model selection process. This level of documentation is critical for reproducibility attempts of the AI models [33, 34], and with this also for the certification process, as it provides clear insights into the design and functionality of the model, facilitating a thorough evaluation. While solid documentation is essential for any certification, it is especially critical for our sample certification, since no company or active development team is managing the project anymore. Although no active development team provides ongoing support, contacting authors and lead developers of the EmoPy framework articles and code, proved beneficial. They kindly addressed questions about the framework and the RIOT setup. Beyond this input, we needed to extract all necessary information from the provided documentation. This reliance on static resources poses a limitation compared to standard certification processes, where ongoing interactions with active developers should be possible.
@@
From a technical perspective, we considered the suitability of the EmoPy framework for sample certification a key factor. Its technical characteristics and comprehensive documentation make it well-suited for the certification process. Additionally, the integration of EmoPy within the RIOT project provides a complete system context, which is essential for certification. Traditional certification catalogues often struggle to validate standalone ML models. However, the RIOT project offers a comprehensive framework where the AI component is embedded within a broader system. This integration is critical, as certification typically applies to an entire system rather than individual components. All these factors influenced our choice to use the EmoPy framework and the RIOT project for our sample certification. Key considerations included its potential relevance to the EU AI Act, its open-source nature ensuring transparency, and its suitability for a sample certification. Additionally, the comprehensive system context that the RIOT project provides reinforced this decision, enabling an effective and thorough certification process. It is also important to note that both EmoPy and RIOT were not originally intended for certification by their creators. However, they serve well for this academic exercise, as the focus is on the certification procedure rather than the system’s quality or real-world applicability.
@@ 4.2 Brief Overview of the EmoPy Framework
The EmoPy framework provides multiple neural network architectures for facial expression recognition, including ConvolutionalNN, TransferLearningNN, and ConvolutionalLstmNN. These architectures vary in complexity, with the ConvolutionalNN beingthe simpleste, and TransferLearningNN (using Google’s Inception-v3) being the most complex. The authors experimented with different architectures and found that the ConvolutionalNN provides the best overall performance [35]. The EmoPy documentation suggests using two publicly available datasets for training and evaluation: the Microsoft FER2013 dataset and the Extended Cohn-Kanade dataset. The key goals of the EmoPy project are to provide free, open-source, and easy-to-use facial expression recognition capabilities, and to advance research in this field by making the models and datasets publicly available. The Framework was also used for the emotion recognition in the RIOT Project.
@@ 4.3 Brief Overview of the RIOT Project
Figure 1: RIOT installation. This image shows the RIOT Art installation in New York City in 2018. One participant is standing in front of the screen taking part in the experience [35].
@@
4.4 The AI Application to be Certified
For this sample certification, the RIOT installation provides the complete system context for the certification process. The EmoPy framework, which provides ML-Models to detect emotions, serves as the core framework within this installation. Our certification process focuses on validating the AI system within the RIOT context. The surrounding code, setup, and information that comprise the entire art installation are relevant, as we cannot certify the AI system independently of these components. However, this approach has shortcomings, as not all the required information is available. Our approach allows for an exploration of the certification process while acknowledging its limitations and academic nature. To facilitate the certification process, we prepared a complete overview of all available information about the AI application, which is summarized below. We also compiled a complete technical overview of the system, that we certified.
4.4.1 Compilation of System Resources
@@
4.4.2 Technical Overview
The AI Application that we attempted to certify has the following specifications: The developers trained the ConvolutionalNN from the EmoPy Framework, as this architecture had the best performance during early testing. The layers and parameters of the ConvolutionalNN can bee seen in Table 1.
Layer Name Type Output Shape Parameters
conv2d_1 Conv2D (None, 61, 61, 10) 170
conv2d_2 Conv2D (None, 58, 58, 10) 1610
max_pooling2d_1 MaxPooling2D (None, 58, 29, 5) 0
conv2d_3 Conv2D (None, 55, 26, 10) 810
conv2d_4 Conv2D (None, 52, 23, 10) 1610
max_pooling2d_2 MaxPooling2D (None, 52, 11, 5) 0
flatten_1 Flatten (None, 2860) 0
dense_1 Dense (None, 4) 11444
Table 1: Summary of model layers and parameters. [35]
In total the model has 15640 trainable parameters and has 4 outputs to be able to classify 4 emotions. The model architecture consists of multiple convolutional and max-pooling layers, which gradually reduce the spatial dimensions while extracting hierarchical features. A final dense layer connects the extracted features to the output layer for classification of the emotions: anger, fear, calm or surprise.
To train the model, two datasets were used. The FER2013 dataset contains over 35,000 facial expression images across 7 emotion classes, while the Cohn-Kanade dataset includes 327 facial expression sequences. However, for this application only the subset of four emotions (anger, fear, calm and surprise) were used. The number of images for each of these emotions can be seen in Table 2. To increase the size and suitability of the training and validation datasets, the authors applied data augmentation techniques. The training process begins by splitting the dataset into training and validation subsets. During training, the network weights are iteratively adjusted to minimize the loss between predicted and labelled emotions. To mitigate overfitting, the authors monitored the gap between training and validation accuracy. This approach prevents overfitting and ensures the model generalizes well by using unseen validation data during training. They measured performance using training and validation accuracy, to help refine the models [31].
Emotion FER2013 Quantity Cohn-Kanade Quantity
Anger 4953 45
Fear 5121 25
Calm 6198 0
Surprise 4002 83
Table 2: Emotion distribution comparison between FER2013 and Cohn-Kanade datasets for used emotion subset [35].
The developers tested the model and also analysed the confusion matrices to gain insights into the model’s performance [35]. Some confusion matrices are also published [31]. The ones relevant to this application with the used emotion subset for both the FER2013 and the Cohn-Kanade Dataset are also published, and for brevity we combined them into one confusion matrix that represents the AI-systems performance for our certification attempt. We considered that the Cohn-Kanade dataset is approximately 100 times smaller than the FER2013 dataset. The resulting matrix can be seen in Figure 2. The data used during training to create this matrix consists only of the discussed datasets (FER2013 and Cohn-Kanade). However, the final model was trained on these datasets with additional data generated through data augmentation.
Figure 2: Confusion matrix of the trained ConvolutionalNN [35].
@@ 5.4 Protection Requirement Analysis
The protection requirement analysis serves as an important first step in the certification process, identifying the risk dimensions that require more in-depth analysis. This analysis involves evaluating the potential impact of the AI system across various dimensions such as fairness, reliability, and data protection. We examined all dimensions and identified several with medium risk. A summary of the Protection Requirement Analysis can be seen in Table 3. For the purposes of this work, we selected two of the required risk dimensions, namely reliability and fairness, for detailed analysis. This selection helps us to focus the sample certification and manage the scope of the certification process. Exploring these two dimensions is also sufficient to understand the certification procedure and draw the appropriate conclusions.
Dimension Risk Category Reasoning
Fairness Medium Processes personal data (facial images), output linked to personal behaviour but has limited impact on rights.
Autonomy and Control Medium Temporarily influences user perception during the interactive film experience.
Transparency Low Lack of transparency does not affect safety or usefulness significantly.
Reliability Medium Misclassification can degrade user experience but does not cause major harm.
Safety and Security Low No risk of physical injury or significant financial damage from malfunction.
Data Protection Medium Uses personal data (live facial images), though not particularly sensitive or secret data.
Table 3: Protection Requirements Analysis Summary
@@ 5.5 Risk Analysis
The risk analysis forms the core of the certification process. We carried out this step by working through a questionnaire from the Fraunhofer Catalogue. The questions addressed different aspects of the selected risk dimensions. For each dimension, we covered topics such as data quality, model design, testing procedures, and operational considerations. A summary of the conclusion of our Risk Analysis for our investigated dimension can be seen in Table 4. Following the individual dimension analyses, we conducted a cross-dimensional assessment to identify potential trade-offs or interactions between the examined dimensions. This step is crucial to ensure a complete understanding of the AI system’s performance and risks. In our sample certification, we were unable to conclude the certification attempt with a positive result, but found significant shortcomings of the AI Application.
Dimension Summary of Risk Analysis
Reliability The system performs well enough within its defined scope, but lacks complete documentation and robustness testing. It is certifiable with improvements in testing and documentation.
Fairness Insufficient analysis of potential biases and discriminatory behaviour. No clear metrics or target groups defined for fairness. Not certifiable without significant further improvements.
Table 4: Certification Summary for Evaluated Dimensions
In the following section, we discuss in detail the challenges we encountered during this certification process. Based on these experiences, we draw conclusions. Additionally, we discuss comparisons with two other certification catalogues to provide a broader perspective on AI certification methodologies.
@@ 7 Conclusion and Future Research Directions
This work-in-progress presents our ongoing research into the practical application of AI certification catalogues, providing insights into both the capabilities and limitations of current certification approaches, as well as the challenges we encountered in certifying an open-source AI system. The implementation of the Fraunhofer AI Assessment Catalogue demonstrated its effectiveness as a comprehensive certification tool, particularly in its systematic approach to evaluate AI systems. We also found that the approach is, at times, bulky and time-consuming in some areas.
In future works, other approaches and certification catalogues should be considered, to potentially streamline the process. The certification process highlights the critical importance of complete system documentation and active engagement from the development team. The absence of these elements can negatively impact the certification process and even make certification infeasible. During the certification attempt, we identified the key characteristics an AI system must possess to be adequately certified. Before selecting a public AI system for a sample certification, auditors should consider factors like thorough documentation and accessibility to the development team. These findings directly address the initial research objectives by identifying both useful aspects of this certification approach and areas for improvement. They also provide practical insights into the certification process and its limitations. Future work in this field could focus on developing more flexible certification methodologies that accommodate various system states and development scenarios.
Acknowledgements: This work was supported by the FFG COMET program, and the strategic COMET project “KnowCERTIFAI” led by Know Center Research GmbH.